ADNAN MUHAMMED

Grade 12,

Zns solubility in water 0.97g/l we mixed 2 moles of zn(N03)2 in 5L solution find new solubility of Zns

Zns solubility in water 0.97g/l we mixed 2 moles of zn(N03)2 in 5L solution find new solubility of Zns

Grade:12th pass

2 Answers

Samyak Jain
5 years ago
Solubility of ZnS in water is 0.97 g/L = (0.97 / 97) mol/L = 0.01 M.
Solubility product Ksp = (0.01 M)(0.01 M) = 10–4 M2.
If we mix 2 moles of Zn(NO3)2 i.e. 2 moles of Zn2+ ions in 5L solution, the added concentration of Zn2+ ions
is 2/5 = 0.4 M.
         ZnS(s)     \rightleftharpoons    Zn2+(aq)    +     S2–(aq)
Conc.                      s + 0.4             s              Ksp = (s + 0.4)(s)
\Rightarrow 10–4 = (s + 0.4)(s)  \Rightarrow  s2 + (0.4) s – 10–4 = 0.
s = [– 0.4 + \sqrt{(0.4)^2 + 4*10^-^4}] / 2  =  [– 0.4 + \sqrt{16*10^-^2 + 4*10^-^4}] / 2
  = [– 0.4 + 10–2 \sqrt{1604}] / 2  =  – 0.2 + 10–2 \sqrt{401}  =  – 0.2 + 0.20025  =  0.00025
s = 2.5 x 10–4 M = 2.5 x 97 x 10–4 g/L = 0.02425 g/L = new solubility of ZnS.
mbappe6789
3 years ago
Solubility of ZnS -0. 01M
Ksp-10-4
Ksp-(s+0.4) (s) 
Ksp-0.4s   (as 0.4is much greater than s... We consider only 0.4) 
10-4 *1/0.4-s
s equal to 2.5*10-4